# On Biologically Plausible Learning in Continuous Time

### v1

@@ 5.2 Layerwise Error Routing
The layerwise error routing topology propagates error from the output layer step by step to previous layers via $V_{l}^{\top}\bm{\epsilon}_{l}$. This corresponds to a KP or weight mirroring scheme where the backward weights $V_{l}$ are learned and converge to transposes of the forward weights. Figure 6 examines the accuracy of the layerwise error routing topology on a synthetic two-dimensional circles dataset. This task is more nonlinear than the $7\times 7$ downsampled MNIST task, and requires deeper ReLU networks to represent an accurate decision boundary.
@@ 5.3 Biological Timescales
A central result of our study is that robust learning requires plasticity windows that outlast the stimulus duration by at least an order of magnitude, placing $\tau_{\mathrm{plas}}$ firmly in the few-second range. This prediction is biologically plausible and experimentally testable, and it narrows the functional range of $\tau_{\mathrm{plas}}$, which has so far been the least constrained timescale in biology.
This requirement emerges when we parameterize our continuous-time networks with biologically motivated constants and evaluate their performance. Real neural circuits operate on nested timescales, from millisecond synaptic conductances to second-scale plasticity windows and slower homeostatic processes. Yet most algorithmic work abstracts away these temporal constraints. By contrast, our simulations demonstrate that the hierarchy $\tau_{\mathrm{prop}}{}\ll\tau_{\mathrm{plas}}\ll\tau_{\mathrm{dec}}$, characteristic of cortical tissue, is sufficient to support effective learning.
@@ 6 Discussion
Our work set out to ask whether biologically plausible learning algorithms can function in continuous time, without relying on artificial phase separation, and at realistic neural timescales. By casting inference and learning as systems of coupled ODEs, we showed that rules such as SGD, FA, DFA, and KP emerge naturally as limiting cases, and that robust learning depends on the temporal overlap of input and error signals. Our simulations further demonstrate that biologically measured timescales are sufficient to support stable learning, leading to a concrete prediction that plasticity windows must persist on the order of seconds.
Limitations. Our empirical evaluation focuses on small datasets and simple tasks due to computational constraints: training of these systems requires explicit forward integration over long time periods of high-dimensional, coupled ODEs whose timescales span over 6 orders of magnitude. This is computationally intensive and numerically delicate (high stiffness, tight tolerances, long horizons), making large benchmarks prohibitive without extensive engineering. Additionally, our neurons are rate-based and simplified, and many biophysical phenomena (spikes, neuromodulatory heterogeneity, etc.) are not modeled. Despite these simplifications, the success of our continuous-time, overlap-driven networks is promising, and future work may address incorporating more faithful biophysics.
Temporal overlap as the operative constraint. In our ODE formulation, the expected update of a synapse is proportional to the temporal cross-correlation between its presynaptic drive $\mathbf{x}$ and a locally projected error $V_{l}^{\top}\bm{\epsilon}_{l}$, low-pass filtered by the plasticity kernel with time constant $\tau_{\mathrm{plas}}$. For piecewise-constant presentations, this yields the analytic overlap kernel $(T-|\Delta|)_{+}$. Failure bands seen at $|\Delta|\approx T$ confirm this prediction: deeper networks accumulate propagation delays that reduce overlap in early layers, degrading learning. This perspective unifies algorithms like FA, DFA, and KP under a single causal principle: a synapse can learn if and only if temporally overlapping input and error reach it. Weight symmetry or fixed feedback are not essential; overlap is.
Plasticity timescale requirement. From the overlap kernel analysis in Section 4, robust learning requires the plasticity kernel to be effectively flat over a presentation window $T$. Allowing at most a fractional attenuation $\eta$ across the window yields $\tau_{\mathrm{plas}}\geq T/\ln\!\big(1/(1-\eta)\big).$ Empirically we observe a transition near $\tau_{\mathrm{plas}}/T\approx 40$, i.e., $\eta\approx 1-e^{-1/40}\approx 0.025$. With $T=50\,\mathrm{ms}$, this places the threshold at $\tau_{\mathrm{plas}}\gtrsim 2\,\mathrm{s}$. Thus, in our protocols, seconds-scale eligibility is sufficient and—absent longer-range tags—apparently necessary; deeper networks and finite propagation tighten this bound by further reducing effective overlap.
Depth and architectural shortcuts. Unlike digital networks, where error gradients are computed simultaneously for all layers, continuous-time networks suffer cumulative delay with depth. This makes learning increasingly fragile as networks grow deeper. Biological circuits appear to mitigate this through architectural shortcuts: skip and feedback pathways are implemented as interareal projections that bypass intermediate layers and, in the neocortex, create shortcuts across the cortical hierarchy (Douglas and Martin, 2004). These connections can provide faster routes for both inference and error signals, preserving temporal overlap in early layers. This interpretation links our results to HSP theory, which argues that arbitrary backward pathways can suffice for credit assignment—our analysis suggests that their role is to preserve overlap, not to enforce symmetry.
Timescales as a design principle. We find that biologically measured timescales are sufficient to support robust learning in our framework. The observed hierarchy $\tau_{\mathrm{prop}}{}\ll\tau_{\mathrm{plas}}\ll\tau_{\mathrm{dec}}$ mirrors the organization of cortical tissue: millisecond synaptic currents, second-scale eligibility traces, and slow homeostatic decay. In simulation, this separation creates a broad regime where learning is stable and accurate, even in the presence of delays and noise. These results suggest that the brain’s hierarchy of clocks may be a functional requirement for maintaining gradient-aligned learning in continuous time.
Implications for hardware. Analog or neuromorphic implementations face the same challenges as biology: finite propagation speeds, integration imperfections, and heterogeneous device dynamics. Our results indicate that hardware does not need to enforce strict weight symmetry, but must preserve overlap between input and error signals across relevant timescales. Designing circuits with fast signal propagation, intermediate eligibility accumulation, and slow decay may therefore be a practical recipe for stable analog learning systems.
Broader perspective. Our continuous-time formulation gives a model for biological and physical learning where inference and plasticity unfold together, constrained by finite propagation and integration dynamics. In this setting, the essential ingredient is not weight transport or symmetry, but the preservation of temporal overlap between inputs and errors within a hierarchy of timescales.
This perspective yields concrete predictions for biology. Deeper circuits are predicted to be more fragile to propagation delays, motivating anatomical shortcuts (skip and feedback pathways) that preserve overlap in early layers. More importantly, robust learning requires plasticity dynamics that outlast the stimulus window by at least an order of magnitude; with cortical integration times of tens of milliseconds, this places the effective eligibility regime in the seconds range. Thus, we predict that eligibility traces supporting error-driven learning must persist at least on the order of seconds—a regime that experimentalists can directly probe by manipulating trace duration or disrupting long-range feedback pathways.
More broadly, our analysis links cortical physiology, algorithmic proposals like FA/DFA/KP, and the practical requirements of analog hardware. It points to a unifying condition—learning succeeds when input and error are temporally correlated at each synapse—that provides both a mechanistic hypothesis for neuroscience and a design rule for neuromorphic systems.
@@ Appendix A Dynamics of a Single Neuron
Appendix B Datasets & Methodology
We use two datasets in our evaluation of these models. The $7\times 7$ downsampled MNIST dataset is the standard MNIST (LeCun, 1998) dataset that has been downsampled with $4\times 4$ average pooling. The circles dataset is the make_circles dataset from scikit-learn (Pedregosa et al., 2011).
[…]
Appendix C Comparison to SGD

### v2

@@ 5.2 Layerwise Error Routing
The layerwise error routing topology propagates error from the output layer step by step to previous layers via $V_{l}\bm{\epsilon}_{l}$. This corresponds to a KP or weight mirroring scheme where the backward weights $V_{l}$ are learned and converge to transposes of the forward weights. Figure 6 examines the accuracy of the layerwise error routing topology on a synthetic two-dimensional circles dataset. This task is more nonlinear than the $7\times 7$ downsampled MNIST task, and requires deeper ReLU networks to represent an accurate decision boundary.
@@ 5.3 Biological Timescales
A central motivation for these algorithms is their relevance to biological learning. Real neural circuits operate on multiple nested timescales, from millisecond synaptic conductances to second-scale plasticity windows and slower homeostatic processes. Yet most algorithmic work has abstracted away these temporal constraints. Here, we parameterize our continuous-time networks with biologically motivated constants and show that they learn effectively on timescales observed in the brain. Of the three time constants relevant in our model, $\tau_{\mathrm{plas}}$ is the least constrained biologically. Our results therefore provide a new theoretical prediction that narrows its plausible functional range.
A central result of our study is that robust learning requires plasticity windows that outlast the stimulus duration by at least an order of magnitude, placing $\tau_{\mathrm{plas}}$ firmly in the few-second range. This prediction is biologically plausible and experimentally testable, and it narrows the functional range of $\tau_{\mathrm{plas}}$, which has so far been the least constrained timescale in biology. Our simulations demonstrate that the hierarchy $\tau_{\mathrm{prop}}{}\ll\tau_{\mathrm{plas}}\ll\tau_{\mathrm{dec}}$, characteristic of cortical tissue, is sufficient to support effective learning.
@@ 6 Discussion
We asked whether feedback-alignment learning can operate in continuous time, without artificial phase separation and at realistic neural timescales. To answer this question, we formulated a coupled ODE model of feedback alignment and characterized its behavior as a function of error delays and plasticity time constants. By embedding neural activity and synaptic plasticity within a single set of differential equations, we show that learning can emerge from continuous-time dynamics without any explicit synchronization.
Our model demonstrates that the central requirement is not weight symmetry, but temporal correlation, which has been an belief in reinforcement learning and neuroscience (Gerstner et al., 2018). Synaptic updates are proportional to the filtered cross-correlation between presynaptic activity and locally projected error signals. When these signals overlap in time, learning approximates gradient descent; when overlap vanishes, updates become biased and performance collapses. In this view, weight transport is not the primary obstacle to biological plausibility. The deeper constraint is maintaining input–error coincidence within the plasticity window of each synapse.
This perspective clarifies the functional role of timescale separation. The hierarchy $\tau_{\mathrm{prop}}{}\ll\tau_{\mathrm{plas}}\ll\tau_{\mathrm{dec}}$} ensures that neural states track inputs rapidly, plasticity integrates over stimulus windows, and long-term decay does not interfere with short-term credit assignment. Timescale separation acts as a structural condition for stable dynamical learning, independent of the specific feedback topology (FA, DFA, or KP).
Therefore, within the limits of a rate-based dynamical model, our results suggest that feedback-alignment–style algorithms can be compatible with biological timescales. These results also suggest that it may be difficult to scale such algorithms to deeper networks in continuous time, though this may not be a major problem because it is yet unclear how “deep” biological brains actually are, and there certainly exist parts of the human brain that can be modeled by shallow networks (Lillicrap et al., 2020; Akrout et al., 2019).
These results could extend beyond biological modeling. Any analog physical substrate must contend with finite propagation delays and heterogeneous device dynamics. Our analysis suggests that such systems need not enforce exact symmetry between forward and backward pathways. Instead, they must preserve sufficient temporal overlap between signals driving synaptic change. Continuous-time learning is therefore feasible in physical systems so long as this overlap condition is met. More broadly, this work connects error-propagation algorithms to the theory of dynamical systems. Learning can be understood as the accumulation of temporally filtered correlations under coupled state–parameter evolution. In this formulation, gradient alignment arises as a consequence of sustained coincidence rather than architectural symmetry. This suggests a unified lens through which to interpret cortical learning, feedback-alignment algorithms, and hardware implementations as instances of the same dynamical principle.
@@ Appendix A Dynamics of a Single Neuron
Appendix B Discrete-time limits
The following discrete-time limits corresponding to SGD, FA, DFA, and KP follow directly from the updates analyzed by Ziyin et al. (2025). In discrete-time FA/DFA, the local modulatory (learning) drive $V_{l}\,\bm{\epsilon}_{l}(t)\in\mathbb{R}^{d_{l}}$ is often written with an additional factor $\sigma_{l}^{\prime}(\mathbf{a}_{l})$ where $\mathbf{a}_{l}=W_{l}\mathbf{z}_{l-1}$. For the ReLU activation used in our experiments, $\sigma_{l}^{\prime}$ is just a binary activity mask, so we treat $V_{l}\,\bm{\epsilon}_{l}$ as the effective gated drive and do not write $\sigma_{l}^{\prime}$ separately. Given the error $\mathbf{e}_{L}$ as the gradient of the loss function with respect to the output layer:
SGD. Let $\mathbf{a}_{l}=W_{l}\mathbf{z}_{l-1}$ and define the backpropagated pre-activation error $\bm{\delta}_{l}:=\partial\mathcal{L}/\partial\mathbf{a}_{l}$. Set the output-layer error source to $\bm{\epsilon}_{L}=\bm{\delta}_{L}$, and for hidden layers set $\bm{\epsilon}_{l}=\bm{\delta}_{l+1}$ with feedback weights $V_{l}=W_{l+1}^{\top}$ (for $l<L$), so that $V_{l}\bm{\epsilon}_{l}=\bm{\delta}_{l}$ (after the implicit ReLU gate). Then $\Delta W_{l}\propto\mathbb{E}\!\left[\bm{\delta}_{l}\,\mathbf{z}_{l-1}^{\top}\right]$, which is exactly the gradient of the loss.
Feedback Alignment (FA). Set $V_{l}$ to fixed random matrices (of shape $d_{l}\times d_{l+1}$) that project the error from layer $l+1$ to layer $l$. Let $\bm{\epsilon}_{L}=\bm{\delta}_{L}$ and $\bm{\epsilon}_{l}=\bm{\delta}_{l+1}$ for $l<L$, so that the local learning drive is $V_{l}\bm{\epsilon}_{l}$. Then $\Delta W_{l}\propto\mathbb{E}\!\left[\big(V_{l}\bm{\epsilon}_{l}\big)\,\mathbf{z}_{l-1}^{\top}\right]$, which is the standard FA update. Gradient alignment arises empirically (Lillicrap et al., 2016).
Direct Feedback Alignment (DFA). Set $V_{l}$ to fixed random matrices (of shape $d_{l}\times d_{L}$). Broadcast the output-layer error to all hidden layers by setting $\bm{\epsilon}_{l}=\bm{\delta}_{L}$ for all $l$. Then the local learning drive is $V_{l}\bm{\delta}_{L}$ and $\Delta W_{l}\propto\mathbb{E}\!\left[\big(V_{l}\bm{\delta}_{L}\big)\,\mathbf{z}_{l-1}^{\top}\right]$, which is the DFA update: each layer learns from a direct projection of the global error (Nøkland, 2016).
Kolen–Pollack (KP) / Weight-Mirror Methods. Allow $V_{l}$ to evolve under the full plasticity rules in equation 1. In the layerwise error-routing topology we take $\bm{\epsilon}_{L}=\bm{\delta}_{L}$ and, for $l<L$, $\bm{\epsilon}_{l}=\bm{\delta}_{l+1}$ so that $V_{l}$ is driven by correlations between the forward drive $W_{l}\mathbf{z}_{l-1}$ and the next-layer error source. In expectation, this update combined with decay drives $V_{l}$ toward a transpose-like mapping that approximates $W_{l+1}^{\top}$, recovering a backprop-like learning drive $V_{l}\bm{\epsilon}_{l}$ without enforcing weight transport (Akrout et al., 2019; Kolen and Pollack, 1994).
Appendix C Datasets & Methodology
We use two datasets in our evaluation of these models. The $7\times 7$ downsampled MNIST dataset is the standard MNIST (LeCun, 1998) dataset that has been downsampled with $4\times 4$ average pooling. The circles dataset is the make_circles dataset from scikit-learn (Pedregosa et al., 2011). It is well established that simple linear classifiers achieve surprisingly high accuracy on MNIST, with only modest improvements from deeper architectures (LeCun et al., 2002). We find a linear softmax regressor on $7\times 7$ downsampled MNIST achieves over 89% test accuracy. In contrast, a logistic regressor on the 2-circle concentric-rings dataset achieves only 75% test accuracy, which corresponds to the class imbalance itself (75% outer ring, 25% inner ring).
$W$ are initialized according Xavier normalization (Glorot and Bengio, 2010). $V$ are initialized to a fixed constant 0.1. Classification decisions are read out from output neurons at the very end of each input sample’s presentation. Evaluation is done on frozen $W$ and $V$ dynamics. Each heatmap data point is the average of 3-5 runs, depending on the experiment.
We integrate all continuous-time dynamics using Diffrax’s (Kidger, 2021) Tsit5 solver (Tsitouras, 2011) (a fifth-order explicit Runge–Kutta method with an embedded fourth-order error estimate) equipped with a PID adaptive step-size controller (rtol $2\times 10^{-3}$, atol $10^{-5}$). This setup allows the solver to take large steps during slowly varying segments of the dynamics while automatically refining steps around rapid transients induced by input switches and error onsets. Although the timescales of our systems span several orders of magnitude, we empirically find that they are not so stiff as to require an implicit method; Tsit5 remains stable and efficient under these tolerances. Compared to a fixed-step forward Euler integrator tuned to resolve the fastest timescale, this adaptive scheme reduces wall-clock time by several orders of magnitude while producing indistinguishable learning curves and dynamics.
Appendix D Comparison to SGD